OpenStudy (anonymous):

integral exponents,simplify; (27p^3q^6r^1)^1/3

OpenStudy (anonymous):

Find the distance between the points (6, 5) and (4, 3).

OpenStudy (anonymous):

?

OpenStudy (kirbykirby):

\[(27p^3q^6r^1)^{1/3}=27^{1/3}(p^3)^{1/3}(q^6)^{1/3} (r^1)^{1/3}\]

OpenStudy (anonymous):

and then i use (ym)n=ymn?

OpenStudy (kirbykirby):

I'm guessing you mean \((y^m)^n=y^{mn}\) , then yes :)

OpenStudy (anonymous):

so the answer will be 27^1/3 p^1 q^2 r^1/3?

OpenStudy (kirbykirby):

yes. Although, you can simplify \(27^{1/3}\) even more!

OpenStudy (anonymous):

okay then 27^1/3 will be 9?

OpenStudy (anonymous):

9 p^1 q^2 r^1/3?

OpenStudy (kirbykirby):

not quite. \(27=3\times 3 \times 3 = 3^3\) So: \(27^{1/3}=(3^3)^{1/3}=3^1=3\)

OpenStudy (kirbykirby):

also, you don't need to write \(p^1\), when you have 1's as exponents, you can just leave it as \(p\)

OpenStudy (anonymous):

okay.thanks. :)

OpenStudy (kirbykirby):

np